OpenStudy (anonymous):

MEDAL!!! Solve for x. 5x + 3(x - 2) = -2(x + 1)

OpenStudy (anonymous):

x = 3 I think

OpenStudy (anonymous):

what would that be in fraction form?

OpenStudy (anonymous):

5x+3x-6=-2x-2 8x+2x=4 10x=4 x=4/10

OpenStudy (anonymous):

4/5 mabey? cause thats not an answer

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

what are the answer choices.. I might know

OpenStudy (anonymous):

4/5 4/3 2/5

OpenStudy (anonymous):

Its 4/5 (:

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

no no its 2/5

OpenStudy (anonymous):

4/10=2/5 when u reduce by 2

OpenStudy (anonymous):

@coolaserman I thought it was 2/5 or 4/5 but I went with 4/5